# Pensieve Grader: An AI-Powered, Ready-to-Use Platform for Effortless Handwritten STEM Grading

### v1

@@ Abstract
Grading handwritten, open-ended responses remains a major bottleneck in large university STEM courses. We introduce Pensieve (https://www.pensieve.co), an AI-assisted grading platform that leverages large language models (LLMs) to transcribe and evaluate student work, providing instructors with rubric-aligned scores, transcriptions, and confidence ratings. Unlike prior tools that focus narrowly on specific tasks like transcription or rubric generation, Pensieve supports the entire grading pipeline—from scanned student submissions to final feedback—within a human-in-the-loop interface.
Pensieve has been deployed in real-world courses at over 20 institutions and has graded more than 300,000 student responses. We present system details and empirical results across four core STEM disciplines: Computer Science, Mathematics, Physics, and Chemistry. Our findings show that Pensieve reduces grading time by an average of 65%, while maintaining a 95.4% agreement rate with instructor-assigned grades for high-confidence predictions.
@@ 1 Introduction
    Partial Credit Assignment: Select rubric items and assign scores granularly, rather than outputting only final grades or free-form justifications.
In this paper, we present Pensieve (https://www.pensieve.co), an LLM-powered grading system designed to meet these needs. Pensieve has been deployed in real-world classrooms, processing over 300,000 student responses across more than 20 institutions. Our system integrates transcription, rubric induction, and rubric-aligned grading within a familiar, human-in-the-loop interface that aligns with existing grading workflows. We evaluate its performance through empirical studies measuring alignment with instructor grading and reductions in grading time.
Our results show that Pensieve can automate a substantial portion of the grading pipeline while preserving grading accuracy and instructor oversight. This work offers a scalable, practical solution for LLM-assisted grading and represents a step toward closing the gap between AI capabilities and instructional needs in large-scale education.
@@ 2 Related Works
We review prior work on rubric-based grading, focusing on systems that leverage large language models (LLMs) across the grading pipeline.
2.1 Image-to-Text Transcription
Accurate transcription of handwritten student work, especially in STEM subjects, remains a major challenge. Traditional OCR systems Smith (2007); Garain and Chaudhuri (2007) often fail to interpret complex mathematical notation due to symbol variety and layout ambiguity. Recent approaches Mok et al. (2024) use multimodal LLMs (e.g., GPT-4V) for image-to-text conversion, achieving strong results on clean inputs but still struggling with subtle distinctions—for example, mistaking $3]{2}$ for $3\sqrt{2}$—which can mislead downstream grading. To mitigate hallucinations, hybrid pipelines (e.g., combining Mathpix and LLMs Kortemeyer et al. (2024)) have been proposed, though they trade off scalability due to longer processing times.
2.2 Rubric Generation and Refinement
In traditional settings, graders iteratively adapt rubrics based on observed error patterns. LLMs offer a promising solution to bootstrap or dynamically revise rubrics. Some systems manually refine rubric items post hoc based on autograding inconsistencies Kortemeyer and Nöhl (2025), while others incorporate human feedback or reinforcement learning to improve rubric clarity and coverage Chu et al. (2024); Li et al. (2025). However, studies show that model-generated rubrics often lack the logical depth and contextual awareness of human-designed ones Wu et al. (2025); Shankar et al. (2024), and iterative rubric self-evaluation by LLMs can introduce bias.
2.3 Rubric Item Selection and Prompt-Based Reasoning
In ALAG, especially for math and physics, assigning partial credit is essential for fair grading Singh et al. (2017). A common approach is to break rubrics into binary items and evaluate each independently. Sonkar et al. Sonkar et al. (2024) follow this method in chemistry grading, improving clarity and consistency. Extensions include ensemble voting and uncertainty flags (e.g., “unsure”) to reflect model confidence Chen and Wan (2025).
Prompt engineering further improves grading quality. Chain-of-thought (CoT) prompts Lee et al. (2024) guide LLMs to reason step by step, while Chen et al. Chen and Wan (2024) introduce scaffolded CoT to explicitly align model reasoning with rubric criteria. Xie et al. Xie et al. (2024) enhance inter-response consistency by prompting over batches of student answers. Retrieval-augmented generation (RAG), as shown by Qiu et al. Qiu et al. (2024), reframes grading as a question-answering task grounded in reference rubrics and solutions.
2.4 Grading Interfaces and User Experience
Despite advances in grading accuracy, many systems lack usable interfaces for non-technical instructors. Tools like BodhiTree-EvalPro Nagakalyani et al. (2025) integrate LLMs but are often limited to structured domains (e.g., code) or rely on unfamiliar UI paradigms.
Pensieve addresses this gap by embedding LLM-assisted grading within a familiar, industry-standard interface. It supports rubric editing, partial credit selection, and instructor review, while providing class-level grading analytics such as mean and standard deviation—making AI-powered grading accessible to real-world classrooms.
@@ 3 System Details
This section provides a detailed overview of Pensieve’s design and workflow. The system builds on familiar grading practices while integrating LLM-powered automation to streamline the grading of open-ended student responses.
Instructors can upload student work scans in bulk, or students may upload their assignments individually. When uploaded in bulk, Pensieve automatically matches scans to students by recognizing handwritten names, reducing manual overhead during setup.
@@
While Pensieve supports autograding for multiple-choice formats, this paper focuses on open-ended responses—such as handwritten solutions, text, or code—where LLMs provide the greatest benefit. We now describe the core AI-assisted components that distinguish Pensieve from traditional grading systems.
@@ 3.1 AI Transcription
For open-ended Text/Code questions, Pensieve first transcribes student responses from scanned images. This transcription step uses OCR and language models to convert handwritten or typed text into machine-readable input.
@@ 3.2 AI Rubric Generation & Calibration
A key differentiator of Pensieve is its ability to generate and refine grading rubrics using LLMs. Each rubric consists of one or more items, with each item assigned a point value and a descriptive label. Rubrics can follow either a subtractive scheme—deducting points for common errors—or an additive scheme—awarding points for correct components. Minimum and maximum total scores can also be specified.
After the initial rubric generation, Pensieve supports a calibration process to fine-tune the rubric and grading behavior based on instructor input. This is particularly useful for open-ended questions, where rubric interpretation can be subjective. For example, instructors may differ in how they interpret vague labels such as “insufficient explanation” (see Figure 2).
To calibrate the AI grader, instructors review a small number of AI-graded examples and correct any mistakes by adjusting the rubric selections. Pensieve then uses these corrections to refine both the rubric interpretation and the grading logic.
@@
Over time, Pensieve also learns from historical rubric usage and calibration patterns to generate better-aligned rubric suggestions for future assignments, significantly reducing instructor workload and promoting grading consistency.
@@ 3.3 AI Grading Confidence
Pensieve assigns each AI-generated grade a confidence level—high, medium, or low. Instructors can use these confidence scores to customize the level of oversight they apply to AI-generated results.
@@
To support this flexible oversight, instructors can configure Pensieve’s behavior based on confidence levels:
@@ 3.4 AI Feedback
In large STEM courses, students often receive limited feedback on open-ended questions due to the grading burden. Pensieve addresses this by allowing instructors to generate individualized feedback after grading.
@@ 3.5 AI Summary
To improve grading transparency and help instructors understand AI decisions, Pensieve generates a concise summary of each student response. These summaries highlight the key reasoning behind the selected rubric items and flag major errors, enabling instructors to verify the AI’s work more efficiently.
@@ 4 Evaluation
We evaluate the system from two key perspectives: (1) the growth in the number of autograding tasks completed using Pensieve, and (2) the amount of grading time saved for instructors.
Figure 3: Cumulative Number of Questions Graded Using Pensieve (Jan–Jun 2025)
@@ 4.1 Usage Patterns and Subject-wise Distribution
Figure 4: Distribution of subjects that used Pensieve
Figure 4 illustrates the subject-wise breakdown of autograded questions. Computer Science dominates, as expected, due to the structured nature of code-based answers and the minimal need for handwriting transcription. More notably, Mathematics and Physics also account for a substantial portion of usage, demonstrating that Pensieve’s LLM-powered transcription and rubric reasoning pipelines can effectively handle symbol-heavy, multi-step problems at the university level. As discussed earlier, these subjects present significant challenges due to ambiguous notation and complex contextual reasoning, yet the system still meaningfully reduces grading effort.
@@ Conclusion
We presented a system that leverages AI to significantly reduce grading time—by an average of 65%—across more than 300,000 open-ended questions from four major college-level STEM domains: Computer Science, Mathematics, Physics, and Chemistry. This reduction is achieved without any meaningful loss in grading accuracy, even for complex, free-form responses.
Pensieve integrates large language models into every stage of the grading workflow, from rubric generation and calibration to transcription, grading, and feedback. Its adaptive learning and instructor-in-the-loop calibration mechanisms ensure that grading aligns with diverse teaching styles while minimizing manual effort.
Designed for practical deployment, Pensieve is readily usable by any instructor. By automating repetitive evaluation tasks, our system enables educators to focus more on high-impact activities such as student interaction and support—while also providing faster, more detailed feedback to learners.

### v2

@@ Abstract
Grading handwritten, open-ended responses remains a major bottleneck in large university STEM courses. We introduce Pensieve Grader (https://www.pensieve.co), an AI-assisted grading platform that leverages large language models (LLMs) to transcribe and evaluate student work, providing instructors with rubric-aligned scores, transcriptions, and confidence ratings. Unlike prior tools that focus narrowly on specific tasks like transcription or rubric generation, Pensieve Grader supports the entire grading pipeline—from scanned student submissions to final feedback—within a human-in-the-loop interface.
Pensieve Grader has been deployed in real-world courses at over 20 institutions and has graded more than 300,000 student responses. We present system details and empirical results across four core STEM disciplines: Computer Science, Mathematics, Physics, and Chemistry. Our findings show that Pensieve Grader reduces grading time by an average of 65% without a meaningful loss in accuracy.
@@ 1 Introduction
    Confidence metric: Let instructors know whether they should skip reviewing AI-generated grade or not based on the expected accuracy.
  4. 4.
    AI Summary & Feedback: Provide a concise summary of any mistakes in the student’s work to help instructors quickly verify the AI-generated grade during grading. After grading, generate detailed feedback for each student to support their learning.
In this paper, we present Pensieve Grader (https://www.pensieve.co), an LLM-powered grading system designed to meet these needs. Pensieve Grader has been deployed in real-world classrooms, processing over 300,000 student responses across more than 20 institutions. Our system integrates transcription, rubric induction, and rubric-aligned grading within a familiar, human-in-the-loop interface that aligns with existing grading workflows. We evaluate its performance through empirical studies measuring alignment with instructor grading and reductions in grading time.
Our results show that Pensieve Grader can automate a substantial portion of the grading pipeline while preserving grading accuracy and instructor oversight. This work offers a scalable, practical solution for LLM-assisted grading and represents a step toward closing the gap between AI capabilities and instructional needs in large-scale education.
@@ 2 Related Works
Numerous studies have investigated the accuracy and reliability of AI-based autograders in STEM education, particularly those utilizing large language models (LLMs). For instance, Qiu et al. (2024) explores applications in biology, while Nagakalyani et al. (2025) focuses on grading computer science assignments. Meanwhile, Liu et al. (2024); Kortemeyer et al. (2024) examines the unique challenges of grading mathematics and physics problems, which often require verbal reasoning and natural language interpretation.
When categorizing prior work based on the techniques employed to improve grading pipelines, prompt engineering frequently emerges as a central strategy. Several studies adopt Chain-of-Thought (CoT) prompting (Lee et al. (2024), Chen and Wan (2024)), while others incorporate retrieval-augmented generation (RAG) Qiu et al. (2024). Despite these advances, many approaches operate under idealized assumptions, such as having access to both pre-defined rubrics and transcribed student responses. For example, Liu et al. (2024) uses a combination of MathPix  [^1] https://mathpix.com and ChatGPT for transcription but presumes the rubric is provided. Similarly, Chu et al. (2024) and Li et al. (2025) iteratively refines rubrics but still assumes textual input from students is available.
To be viable in real university courses, an AI grading system must handle cases where the model fails to produce accurate results. For example, Kortemeyer and Nöhl (2025) leverage model confidence scores to flag grades that require human review. Another promising enhancement is the inclusion of AI-generated summaries that help instructors quickly identify student mistakes. However, few prior works have explored this aspect—likely because their systems lack mechanisms for confidence differentiation or integrated human-in-the-loop review pipelines. In addition, delivering timely feedback is a crucial aspect of the grading process. While some studies have explored using large language models (LLMs) to generate feedback, they often lack alignment with grading rubrics or rely on vague, qualitative criteria (e.g., "Code Readability"). To the best of our knowledge, Pensieve Grader is the first system to offer a comprehensive, end-to-end pipeline that includes raw image transcription, rubric-based evaluation, confidence-driven review prioritization, and structured summary and feedback generation.
@@ 3 System Details
This section provides a detailed overview of Pensieve Grader’s design and workflow. The system builds on familiar grading practices while integrating LLM-powered automation to streamline the grading of open-ended student responses.
Instructors can upload student work scans in bulk, or students may upload their assignments individually. When uploaded in bulk, Pensieve Grader automatically matches scans to students by recognizing handwritten names, reducing manual overhead during setup.
@@
While Pensieve Grader supports autograding for multiple-choice formats, this paper focuses on open-ended responses—such as handwritten solutions, text, or code—where LLMs provide the greatest benefit. We now describe the core AI-assisted components that distinguish Pensieve Grader from traditional grading systems.
@@ 3.1 AI Transcription
For open-ended Text/Code questions, Pensieve Grader first transcribes student responses from scanned images. This transcription step uses OCR and language models to convert handwritten or typed text into machine-readable input.
@@ 3.2 AI Rubric Generation & Calibration
A key differentiator of Pensieve Grader is its ability to generate and refine grading rubrics using LLMs. Each rubric consists of one or more items, with each item assigned a point value and a descriptive label. Rubrics can follow either a subtractive scheme—deducting points for common errors—or an additive scheme—awarding points for correct components. Minimum and maximum total scores can also be specified.
After the initial rubric generation, Pensieve Grader supports a calibration process to fine-tune the rubric and grading behavior based on instructor input. This is particularly useful for open-ended questions, where rubric interpretation can be subjective. For example, instructors may differ in how they interpret vague labels such as “insufficient explanation” (see Figure 2).
To calibrate the AI grader, instructors review a small number of AI-graded examples and correct any mistakes by adjusting the rubric selections. Pensieve Grader then uses these corrections to refine both the rubric interpretation and the grading logic.
@@
Over time, Pensieve Grader also learns from historical rubric usage and calibration patterns to generate better-aligned rubric suggestions for future assignments, significantly reducing instructor workload and promoting grading consistency.
@@ 3.3 AI Grading Confidence
Pensieve Grader assigns each AI-generated grade a confidence level—high, medium, or low. Instructors can use these confidence scores to customize the level of oversight they apply to AI-generated results.
@@
To support this flexible oversight, instructors can configure Pensieve Grader’s behavior based on confidence levels:
@@ 3.4 AI Feedback
In large STEM courses, students often receive limited feedback on open-ended questions due to the grading burden. Pensieve Grader addresses this by allowing instructors to generate individualized feedback after grading.
@@ 3.5 AI Summary
To improve grading transparency and help instructors understand AI decisions, Pensieve Grader generates a concise summary of each student response. These summaries highlight the key reasoning behind the selected rubric items and flag major errors, enabling instructors to verify the AI’s work more efficiently.
@@ 4 Evaluation
We evaluate the system from two key perspectives: (1) the growth in the number of autograding tasks completed using Pensieve Grader, and (2) the amount of grading time saved for instructors.
Figure 3: Cumulative Number of Questions Graded Using Pensieve Grader (Jan–Jun 2025)
@@ 4.1 Usage Patterns and Subject-wise Distribution
Figure 4: Distribution of subjects that used Pensieve Grader
Figure 4 illustrates the subject-wise breakdown of autograded questions. Computer Science dominates, as expected, due to the structured nature of code-based answers and the minimal need for handwriting transcription. More notably, Mathematics and Physics also account for a substantial portion of usage, demonstrating that Pensieve Grader’s LLM-powered transcription and rubric reasoning pipelines can effectively handle symbol-heavy, multi-step problems at the university level. As discussed earlier, these subjects present significant challenges due to ambiguous notation and complex contextual reasoning, yet the system still meaningfully reduces grading effort.
@@ Conclusion
We presented a system that leverages AI to significantly reduce grading time—by an average of 65% compared to grading with traditional grading software—across more than 300,000 open-ended questions from four major college-level STEM domains: Computer Science, Mathematics, Physics, and Chemistry. This reduction is achieved without any meaningful loss in grading accuracy, even for complex, free-form responses.
Pensieve Grader integrates LLMs into every stage of the grading workflow, from rubric generation and calibration to transcription, grading, and feedback. Its adaptive learning and instructor-in-the-loop calibration mechanisms ensure that grading aligns with diverse teaching styles while minimizing manual effort.
Designed for practical deployment, Pensieve Grader is readily usable by any instructor. By automating repetitive evaluation tasks, our system enables educators to focus more on high-impact activities such as student interaction and support—while also providing faster, more detailed feedback to learners.
Acknowledgement
We thank Tony Dear, a senior lecturer in Columbia University, for sharing the question and the rubric for Figure 1.